Not sure of order of operation......
f=1
f=45-4e+1
f=45-4+1
does f=42 or does f=40

Answer by checkley71(8403)

F=45-4+1=45+1-4=46-4=42
F=45-4+1=41+1=42 NO MATTER WHICH OPERATION YOU DO FIRST.
THE SEQUENCE OF OPERATIONS IS RESOLVE ALL ITMS WITHIN THE ( ).
THEN pOWERS, MULTIPLY/DIVIDE THEN ADD/SUBTRACT.